OpenStudy (anonymous):

Solve the inequality algebraically. Show your work(a number line) (x+7)(x-4) / x-1 >= 0 Is the answer: 4 >= x >= -7???

myininaya (myininaya):

Is your thingy saying: \[\frac{(x+7)(x-4)}{x} -1 \ge 0\]

OpenStudy (anonymous):

No x-1 is the denominator.

myininaya (myininaya):

\[\frac{(x+7)(x-4)}{x-1} \ge 0\] check around numbers -7, 4 ,1

myininaya (myininaya):

top=0 bottom=0 when this happens you check around the numbers that satisfy those equations